OpenStudy (anonymous):

PLEASE HELP GIVING MEDAL What is the direct linear variation equation for the relationship? y varies directly with x and y = 4 when x = 12. A. y = x + 8 B. 1 y= ---x 3 C. y = x – 8 D. y = 3x

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

can you please help

OpenStudy (anonymous):

@mavagirl552000

OpenStudy (anonymous):

@jigglypuff314

Directrix (directrix):

y = kx 4 = k*12 k = 4/12 k = 1/3 y = (1/3) * x

Directrix (directrix):

k is called the constant of variation http://www.regentsprep.org/regents/math/algebra/AO4/Ldirect.htm

OpenStudy (anonymous):

so D?

Directrix (directrix):

No.

Directrix (directrix):

y = (1/3) * x That is a 1/3, not 3. Try again.

OpenStudy (anonymous):

ohh ok hold up

OpenStudy (anonymous):

B?

Directrix (directrix):

@ACU17 ^^^^^

OpenStudy (anonymous):

so am I correct @Directrix

OpenStudy (anonymous):

?

Directrix (directrix):

>>>> B? Correct

OpenStudy (anonymous):

thanks